OpenStudy (anonymous):

Graph the line 4/3 x how to find points?

OpenStudy (anonymous):

Graph y = 4/3x first.

OpenStudy (danjs):

To graph a line, you need two points. y = 4/3 * x Let x = 0, and x = 1, find y for each. (0,y) and (1,y)

OpenStudy (anonymous):

how do i do that?

OpenStudy (danjs):

If x=0, y = 4/3 * 0 = 0 The first point is x=0, y=0 (x,y) = (0,0)

OpenStudy (danjs):

if x=1, y=(4/3)* 1 = 4/3 (x,y) = (1, 4/3) there you have 2 points, plot them and draw a line through both points.

OpenStudy (anonymous):

Thanks. but how about problems like y=-x?

OpenStudy (danjs):

It is the same for any linear equation. You just need to find 2 points to graph a line. For y=-x For x=0, y=0 (0,0) For x=1, y=-1 (1,-1)

OpenStudy (anonymous):

thanks